how do you find the equation of a line that has the slope and passes through the given point?
m = 2; (3,2)

Answer by tarungehlot(9)

such type of problem are solved by point slope form
in which a slope and a point is given in problem
y- y1 = m( x-x1)
y - 3= 2(x-2)
y-3= 2x -4
y= 2x -1